# On learning racing policies with reinforcement learning

### v1

@@ I Introduction
Ensuring the reliable performance of fully autonomous vehicles in critical corner cases is essential for achieving their promised safety and enabling large-scale deployment. In dynamic scenarios, even small control inaccuracies can lead to crashes. As a result, control strategies must be both robust to uncertainties and capable of operating at the vehicle’s limit to minimize the risk of accidents.
@@
One promising approach to developing robust control algorithms is using Reinforcement Learning (RL). Unlike optimization-based controllers, RL can leverage domain randomization to enhance robustness against modeling errors and effectively handle sparse and discontinuous objectives. Moreover, RL is not constrained by a finite prediction horizon, enabling it to optimize long-term performance beyond the capabilities of traditional optimization-based methods. Previous studies have demonstrated that RL can surpass professional human drivers in car racing simulations [2, 3, 4] and even outperform expert pilots in real-world drone racing [5]. However, despite its success in racing simulations, RL policies have struggled to transfer effectively to real-world scenarios. Consequently, optimization-based methods have consistently outperformed RL policies in real-world racing [6, 7].
In this work, we propose a novel RL-based approach to autonomous racing that not only outperforms state-of-the-art MPC methods but also, to the best of our knowledge, represents the first instance of a RL policy that surpasses an expert human driver in RC racing. Notably, our benchmark human driver achieved fifth place in the Polish Indoor RC Championship, validating the competitive standard against which our autonomous system was evaluated. We systematically explore critical design choices, including domain randomization, actuator dynamics modeling, and policy architecture, which are essential for improving performance, safety, and transferability. The RL policies are trained exclusively in simulation and subsequently validated on the F1TENTH [8] platform under real-world conditions without additional training. Supplementary video materials from the experimental evaluation can be found at https://grzegorzczput.github.io/rl-racing/.
@@
[…]
@@ II-A Model Predictive Control
One of the most popular approaches to designing a controller for autonomous car racing is to leverage MPC algorithms [9, 10, 11]. These methods have been proven effective in real-world deployments. However, a major drawback of optimization-based approaches is their limited control frequency, which is constrained by model complexity and horizon length. This limitation forces users to adopt relatively simple models and significantly shorter horizons than a full lap of a racing track.
Additionally, to ensure fast optimization convergence, the objective function typically includes auxiliary regularization terms [12, 13] or requires simplifications, such as restricting costs to quadratic penalties relative to a predefined trajectory [14]. As a result, the objective function often serves as a proxy task rather than directly optimizing for the controller’s ultimate objective.
@@ II-B Reinforcement Learning in Autonomous Racing
On the other hand, RL offers a promising solution to overcome these limitations. Reinforcement learning has been successfully applied to complex robotic systems such as quadrupeds [15, 16] and drones [17, 5], demonstrating its capability to handle high-dimensional models and long-horizon planning. Unlike MPC, where the computational burden occurs during inference, RL shifts the computational effort to the training phase, enabling efficient real-time execution once training is complete. Moreover, RL is well suited for optimizing tasks with sparse and delayed rewards, such as minimizing lap times in drone racing [18].
The application of RL in autonomous car racing has been studied primarily in simulation [19, 20] and has even been shown to outperform professional human simulator drivers [4]. Nevertheless, implementing end-to-end RL policies in real-world racing remains a non-trivial task. As demonstrated in [6], researchers could not directly train a high-performance policy without the foundation of a geometric controller. Even with this foundation, the resulting policy still failed to achieve performance comparable to model-based approaches such as MPC. Similarly, despite employing real-world policy fine-tuning, the authors [7] could not surpass the performance benchmarks set by optimization-based controllers.
@@ III-A Problem definition
In this paper, we consider the problem of real-world autonomous racing with an F1TENTH car. In general, this is an extremely complex problem that typically requires the integration of multiple subsystems responsible for aggregating the data from sensors, performing localization, opponent detection, planning, and control. However, in this paper, we focus entirely on the control problem in the time-trial scenario, assuming perfect knowledge of the vehicle state. The goal of this task is to find a controller that, based on the knowledge of the vehicle’s state and the racetrack, is able to generate control signals, allowing the car to maximize the progress along the track centerline while remaining within track boundaries. This objective may be considered as an accurate dense approximation of the sparse minimum lap time goal.
@@ III-B Proposed solution
To address the problem of autonomous racing in a time-trial scenario, we begin with developing an accurate simulation of the F1TENTH car. Our goal is to precisely model all relevant aspects of vehicle dynamics, including actuator dynamics and tire characteristics that are lacking in other simulators [21]. We identify system parameters using a long-horizon prediction loss, which helps stabilize the learned models.
@@ III-C Hardware setup
  • Wheel Speed Control Mode: The control input is defined as $a=[\delta_{\text{ref}},\omega_{\text{ref}}]^{T}$, where $\delta_{\text{ref}}$ is the reference steering angle which is directly sent to the servo for steering actuation, and $\omega_{\text{ref}}$ is the reference wheel speed, which is sent to the Electronic Speed Controllers (ESC). The ESC utilizes an internal PID controller to regulate the vehicle’s wheel speed accordingly.
@@
These two control modes are employed due to the differing requirements of human drivers and control algorithms. Human drivers naturally regulate the vehicle by modulating the throttle signal, which in electric vehicles is commonly implemented as motor current control, while autonomous control systems favor wheel speed regulation using high-frequency, low-level PID controllers inside ESC.
To enable both system identification and policy deployment, we employ an OptiTrack motion capture system. OptiTrack provides precise real-time measurements of the vehicle’s position and orientation, which are further processed to estimate velocity in the vehicle frame. The velocity is computed by differentiating position and orientation data using a Savitzky-Golay filter [22].
All training and real-world experiments were conducted on a single L-shaped track, measuring 17 meters in length and 1 meter in width. Despite being constrained by the OptiTrack tracking area, this layout incorporates both a high-speed section and a tight technical sector, providing a balanced environment for evaluating controllers’ performance. This challenging setup ensures a comprehensive assessment of reinforcement learning policies and traditional control algorithms.
@@ III-D Simulation environment
The simulation environment is a key aspect of the reinforcement learning setup, especially for high-performance racing policies. In this paper, we focus solely on the accurate and fast simulation of racing car dynamics, completely neglecting the visual aspect of the simulation, as it will be used only for policy training. Therefore, we want the dynamics model to be (i) fast to evaluate, (ii) difficult or near impossible to exploit, (iii) accurate enough to enable racing at the limit of grip, and finally (iv) close to the hardware setup we use. Such qualities are beyond the scope of general-purpose simulators but can be obtained by using an analytical physics-derived model optimized to be consistent and accurate in long-horizon simulation.
In this paper, we used a single track vehicle dynamics model with the MF6.1 tire model [23]. Moroever, to account for the dynamics of the physical actuators, we model them as a first-order system, identifying the relationship between the reference input and the measured position and velocity by
@@
where $T_{\delta}$ and $T_{\omega}$ are the time constants $T_{\delta}$ for steering actuator and for approximation of PID regulators inside ESC, respectively.
We identified the parameters of the introduced model using the dataset collected with the physical setup introduced in Section III-C. This dataset was collected by an expert human driving on different tracks, with varying ranges of slip angle and slip ratio, trying to cover the entire envelope of the vehicle state space operation region. During data collection, we captured relevant data about the car’s internal state, such as motor RPM and steering angle, as well as information about its position and orientation using the OptiTrack motion capture system. All of these data were collected synchronously at 100Hz. In total, we recorded 36 minutes of driving.
Finally, similarly to [24], we formulated the Mean Squared Error (MSE) loss between the state trajectory roll-out obtained using the dynamics model and ground-truth data on sequences of 85 samples. To minimize this loss, we used stochastic gradient decent optimization, in particular AdamW [25], and backpropagated the loss gradient with respect to the dynamics model’s parameters through time.
@@ III-E Reinforcement Learning Training Procedure
Leveraging the accurate simulation environment detailed in Section III-D, we train our RL policy using the Proximal Policy Optimization (PPO) algorithm [26, 27], in a parallelized setup with 400 simulated instances running concurrently. Each rollout consists of 1024 steps, with a time step of 0.05s. The training was performed using a batch size of 1024. If the agent exceeds track limits during a rollout, it is reset to a random position on the track to ensure full track exploration. During training, the parameters of the dynamics model are randomized at the beginning of each episode by applying multiplicative perturbations with Gaussian noise. Specifically, each parameter $\theta$ is scaled by a factor drawn from $\mathcal{N}(1,\sigma_{dr})$, where $\sigma_{dr}$ is specified in the experimental configuration.
@@
The sum of discounted rewards is calculated with a discount factor of $\gamma=0.99$. The learning rate is decayed from $1\times 10^{-3}$ to $1\times 10^{-4}$ over the course of training. The total training process spans 120 million environment steps, enabling the agent to experience a wide range of driving scenarios and refine its policy accordingly.
Both the actor and critic networks are implemented as Multilayer Perceptronss (MLPs). The actor network consists of two hidden layers, each with 256 neurons, while the critic network comprises two hidden layers with 512 neurons each. Both networks employ the Leaky Rectified Linear Unit (Leaky ReLU) activation function with a negative slope of 0.2. The final output of the actor network is two-dimensional, with each element constrained to the interval $[-1,1]$. These outputs are subsequently scaled by a factor of 0.5 for $\delta_{\text{ref}}$ and 5 for $\dot{\omega}_{\text{ref}}$ to yield the action vector. The steering scaling factor was chosen because 0.5 rad is the maximum limit of the steering actuator. The wheel acceleration limit was set to 5 $m/s^{2}$, as this was the highest longitudinal acceleration measured from the identification dataset.
$a=[\delta_{\text{ref}},\dot{\omega}_{\text{ref}}]^{T}.$ (2)
Here, the reference steering angle $\delta_{\text{ref}}$ is directly applied to the servo for steering actuation, while the reference wheel speed derivative $\dot{\omega}_{\text{ref}}$ is integrated over time, analogously to [28], before being passed to the wheel speed controller. This formulation of speed control mitigates abrupt velocity changes, thereby enhancing policy transferability to the real world.
The reward function, designed to quantify the vehicle’s progress along the track while penalizing boundary violations, is formulated as follows:
where $\mathfrak{s}_{t}$ denotes the progress along the centerline expressed in Frenet coordinates. This reward formulation, or similar variants, is commonly employed in other RL racing problems [3, 4, 5].
The observation configuration of both the actor and critic network is defined as follows:
@@
1. Linear velocities in vehicle frame: longitudinal $v_{x}$ and lateral $v_{y}$. 2. Yaw rate: $\mathfrak{r}$. 3. The angle between the car’s heading and the track heading, expressed in Frenet coordinates: $u$. 4. Lateral distance from the centerline, in Frenet coordinates: $n$. 5. Measured steering angle $\delta$ and the previous control input $\delta_{\text{ref}}$. 6. Measured wheel speed $\omega$, commanded wheel speed $\omega_{\text{ref}}$, and the last control input $\dot{\omega}_{\text{ref}}$. 7. Track information, represented by $N$ points, with curvature $\boldsymbol{c}\in\mathbb{R}^{N}$ and width $\boldsymbol{w}\in\mathbb{R}^{N}$. These $N$ points are sampled uniformly in front of the vehicle at 30-centimeter intervals.
Thus, the observation vector is given by:
@@ IV Experiments
To determine the key factors that contribute to learning RL policies for autonomous racing, we conducted a series of experiments. A major focus of our study was addressing the sim-to-real gap, for which we evaluated various domain randomization strategies (see Section IV-A) and the impact of actuator modeling (see Section IV-D) to identify the most effective approach for improving real-world transferability.
@@ IV-A Domain randomization
[…]
Policy performance was assessed by comparing both lap times and safety, which was quantified by the time-integral of the off-track distance, averaged per lap, measured in $[m\cdot s]$. The off-track distance is defined as
@@
Figure 3: Performance of networks trained with randomized friction and single track parameters. Top: Lap times achieved with different types and $\sigma_{dr}$ of randomization. Bottom: $E_{off}$ and percentage of laps with crashes for different randomization configurations.
The results of the experiment with randomized friction levels and single track parameters are presented in Figure 3. As friction randomization increases, lap times tend to increase while the integrated off-track error decreases. A similar effect was observed when randomizing all single-track parameters; however, overall, single-track parameter randomization caused worse performance, resulting in longer lap times and higher $E_{\text{off}}$ compared to policies trained with only friction randomization. This suggests that under more significant parameter variability, the policy adopts a more conservative driving strategy to remain valid across the entire distribution of dynamic models encountered during training. While domain randomization enhances robustness, excessive perturbations across all parameters can degrade overall performance. When the policy is trained over a wide distribution of models, it is forced to learn a control strategy that generalizes across the entire range of dynamics rather than optimizing for a specific one. Since the policy has no access to action history or recurrent memory to infer the underlying system dynamics at a given moment, it cannot adapt its behavior to a particular model instance. Instead, it must adopt a more cautious, suboptimal racing strategy that remains feasible across all possible parameter variations, ultimately leading to longer lap times.
These findings highlight the necessity of balancing domain randomization to ensure both safety and real-world transferability. Even with precise system identification, a moderate level of randomization is beneficial for mitigating simulation-to-reality discrepancies and improving policy robustness. However, while limited domain randomization enhances generalization, high-performance policies still require accurate system identification, as randomization alone cannot fully compensate for an inaccurate dynamics model. For subsequent experiments, we selected the policy trained with friction randomization at $\sigma_{dr}=0.02$, and we reference it as the base policy. We chose this policy as it offers an acceptable trade-off between speed and robustness.

### v2

@@ I Introduction
Ensuring the reliable performance of fully autonomous vehicles in critical corner cases is essential to achieve their promised safety and enable large-scale deployment. In dynamic scenarios, even small control inaccuracies can lead to crashes. As a result, control strategies must be both robust to uncertainties and capable of operating at the vehicle’s limit to minimize the risk of accidents.
@@
One promising approach to developing robust control algorithms is the use of Reinforcement Learning (RL). Unlike optimization-based controllers, RL can leverage domain randomization to improve robustness against modeling errors and effectively handle sparse and discontinuous objectives. Moreover, RL is not constrained by a finite prediction horizon, allowing it to optimize long-term performance beyond the capabilities of traditional optimization-based methods. Previous studies have demonstrated that RL can surpass professional human drivers in car racing simulations [2][3][4] and even outperform expert pilots in real-world drone racing [5]. However, despite its success in racing simulations, RL policies have struggled to transfer effectively to real-world scenarios. Consequently, optimization-based methods have consistently outperformed RL policies in real-world car racing [6, 7].
In this work, we propose a novel RL-based approach to autonomous racing that not only outperforms state-of-the-art MPC methods but also, to the best of our knowledge, represents the first instance of a RL policy that surpasses an expert human driver in RC racing. Notably, our benchmark human driver achieved fifth place in the Polish Indoor RC Championship, validating the competitive standard against which our autonomous system was evaluated. We systematically explore critical design choices, including domain randomization, actuator dynamics modeling, and policy architecture, which are essential to improve performance, safety, and transferability. The RL policies are trained exclusively in simulation and subsequently validated on the F1TENTH platform [8] under real-world conditions without additional training. Supplementary video materials from the experimental evaluation can be found at https://grzegorzczput.github.io/rl-racing/.
@@
    We evaluated our approach against the state-of-the-art MPC and an expert human RC driver in the real world F1TENTH racing, demonstrating a fastest lap time improvement of $0.078\text{\,}\mathrm{s}$ over MPC and $0.286\text{\,}\mathrm{s}$ over the expert human driver on a challenging $17\text{\,}\mathrm{m}$ racetrack (see Fig. 1).
@@ II-A Model Predictive Control
One of the most popular approaches to designing a controller for autonomous car racing is to leverage MPC algorithms [9, 10, 11]. These methods have proven to be effective in real-world deployments. However, a major drawback of optimization-based approaches is their limited control frequency, which is constrained by model complexity and horizon length. This limitation forces users to adopt relatively simple models and significantly shorter planning horizons than a full lap of a racing track.
Additionally, to ensure fast optimization convergence, the objective function typically includes auxiliary regularization terms [11, 12] or requires simplifications, such as restricting costs to quadratic penalties relative to a predefined trajectory [13]. As a result, the objective function often acts as a proxy, rather than directly optimizing for the controller’s true objective.
@@ II-B Reinforcement Learning in Autonomous Racing
On the other hand, RL offers a promising solution to overcome these limitations. Reinforcement learning has been successfully applied to complex robotic systems such as quadrupeds [14, 15] and drones [16, 5], demonstrating its capability to handle high-dimensional models and long-horizon planning. Unlike MPC, where the computational burden occurs during inference, RL shifts the computational effort to the training phase, enabling efficient real-time execution once training is complete. Moreover, RL is well-suited for optimizing tasks with sparse and delayed rewards, such as minimizing lap times in drone racing [17].
The application of RL in autonomous car racing has been studied primarily in simulation [18, 19] and has even been shown to outperform professional human simulator drivers [4]. Nevertheless, implementing end-to-end RL policies in real-world racing remains a non-trivial task. As demonstrated in [6], researchers could not directly train a high-performance policy without the foundation of a geometric controller. Even with this foundation, the resulting policy still failed to achieve performance comparable to model-based approaches such as MPC. Similarly, despite employing real-world policy fine-tuning, the authors [7] were unable to exceed the performance benchmarks set by optimization-based controllers.
@@ III-A Problem definition
In this paper, we consider the problem of real-world autonomous racing with an F1TENTH car. In general, this is an extremely complex problem that typically requires the integration of multiple subsystems responsible for aggregating the data from sensors, performing localization, opponent detection, planning, and control. However, in this paper, we focus solely on the control problem in a time-trial scenario, assuming perfect knowledge of the vehicle state. The goal of this task is to find a controller that, based on the knowledge of the vehicle’s state and the racetrack, is able to generate control signals, allowing the car to maximize the progress along the track centerline while remaining within track boundaries. This objective may be considered as an accurate dense approximation of the sparse minimum lap time goal.
@@ III-B Proposed solution
To address the problem of autonomous racing in the time-trial scenario, we begin by developing an accurate simulation of the F1TENTH car. Our goal is to precisely model all relevant aspects of vehicle dynamics, including actuator dynamics and tire characteristics that are lacking in other simulators [20]. We identify system parameters using a long-horizon prediction loss, which helps stabilize the learned models.
@@ III-C Hardware setup
  • Wheel Speed Control Mode: The control input is defined as $a=[\delta_{\text{ref}},\omega_{\text{ref}}]^{T}$, where $\delta_{\text{ref}}$ is the reference steering angle sent directly to the servo for steering actuation, and $\omega_{\text{ref}}$ is the reference wheel speed, which is sent to the Electronic Speed Controllers (ESC). The ESC utilizes an internal PID controller to regulate the vehicle’s wheel speed accordingly.
@@
The two control modes are designed to meet the distinct requirements of human drivers and autonomous control algorithms. Human drivers naturally regulate the vehicle by modulating the throttle signal, which in electric vehicles is typically implemented as motor current control. In contrast, autonomous control systems favor wheel speed regulation using high-frequency, low-level PID controllers inside ESC.
To enable both system identification and policy deployment, we employ an OptiTrack motion capture system. OptiTrack provides precise real-time measurements of the vehicle’s position and orientation, which are then processed to estimate velocity in the vehicle frame. The velocity is computed by differentiating the position and orientation data using a Savitzky-Golay filter [21].
All training and real-world experiments were conducted on a single L-shaped track (see Fig. 1), measuring 17 meters in length and 1 meter in width. Despite being constrained by the OptiTrack tracking area, this layout incorporates both a high-speed section and a tight technical sector, providing a balanced environment for evaluating controllers’ performance. This challenging setup ensures a comprehensive assessment of reinforcement learning policies and traditional control algorithms.
@@ III-D Simulation environment
The simulation environment is a key aspect of the reinforcement learning setup, especially for high-performance racing policies. In this paper, we focus solely on the accurate and fast simulation of racing car dynamics, completely neglecting the visual aspect of the simulation, as it will only be used for policy training. Therefore, we want the dynamics model to be (i) fast to evaluate, (ii) difficult or near impossible to exploit, (iii) accurate enough to enable racing at the limit of grip, and finally (iv) close to the hardware setup we use. Such qualities are beyond the scope of general-purpose simulators but can be obtained by using an analytical physics-derived model optimized to be consistent and accurate in long-horizon simulation.
In this paper, we used a single track vehicle dynamics model with the MF6.1 tire model [22], which extends the traditional Pacejka formulation [23] commonly used in autonomous racing [24] by jointly considering lateral and longitudinal tire slip to calculate tire forces. Moreover, to account for the dynamics of the physical actuators, we model them as a first-order system, identifying the relationship between the reference input and the measured position and velocity by
@@
where $T_{\delta}$ and $T_{\omega}$ are the time constants $T_{\delta}$ for the steering actuator and the approximation of the PID regulators inside ESC, respectively.
We identified the parameters of the introduced model using the dataset collected with the physical setup introduced in Section III-C. This dataset was collected by an expert human driver operating across various tracks, intentionally varying slip angles and slip ratios to cover the full operational envelope of the vehicle’s state space. During data collection, we captured relevant data about the car’s internal state, such as motor rotational speed and steering angle, as well as information about its position and orientation using the OptiTrack motion capture system. All of the data were collected synchronously at 100Hz. In total, we recorded 36 minutes of driving.
Finally, similarly to [25], we formulated the Mean Squared Error (MSE) loss between the state trajectory roll-out obtained using the dynamics model and ground-truth data on sequences of 85 samples. To minimize this loss, we used stochastic gradient descent optimization, in particular AdamW [26], and backpropagated the loss gradient with respect to the dynamics model’s parameters through time.
@@ III-E Reinforcement Learning Training Procedure
Leveraging the accurate simulation environment detailed in Section III-D, we train our RL policy using the Proximal Policy Optimization (PPO) algorithm [27, 28], in a parallelized setup with 400 simulated instances running concurrently. Each rollout consists of 1024 steps, with a time step of 0.05s. The training was performed using a batch size of 1024. If the agent exceeds track limits during a rollout, it is reset to a random position on the track’s centerline to enhance the exploration. During training, the parameters of the dynamics model are randomized at the beginning of each episode by applying multiplicative perturbations with Gaussian noise. Specifically, each parameter $\theta$ is scaled by a factor drawn from $\mathcal{N}(1,\sigma_{dr})$, where $\sigma_{dr}$ is specified in the experimental configuration.
@@
The sum of discounted rewards is calculated with a discount factor of $\gamma=0.99$. The learning rate decays from $1\times 10^{-3}$ to $1\times 10^{-4}$ over the course of training. The total training process spans 120 million environment steps, allowing the agent to experience a wide range of driving scenarios and refine its policy accordingly.
Both the actor and critic networks are implemented as multilayer perceptrons. The actor network consists of two hidden layers, each with 256 neurons, while the critic network comprises two hidden layers with 512 neurons each. Both networks employ the Leaky Rectified Linear Unit activation function with a negative slope of 0.2.
@@
The observation configuration for both networks is defined as follows: 1. Linear velocities in vehicle frame: longitudinal $v_{x}$ and lateral $v_{y}$. 2. Yaw rate: $\mathfrak{r}$. 3. The angle between the car’s heading and the track heading, expressed in Frenet coordinates $u$. 4. Lateral distance from the centerline, in Frenet coordinates: $n$. 5. Measured steering angle $\delta$ and the previous control input $\delta_{\text{ref}}$. 6. Measured wheel speed $\omega$, commanded wheel speed $\omega_{\text{ref}}$, and the last control input $\dot{\omega}_{\text{ref}}$. 7. Track information, represented by $N$ points, with curvature $\boldsymbol{c}\in\mathbb{R}^{N}$ and width $\boldsymbol{w}\in\mathbb{R}^{N}$. These $N$ points are sampled uniformly in front of the vehicle at 30-centimeter intervals. Thus, the observation vector is given by:
@@
The final output of the actor network is a two-dimensional vector $\delta_{\text{ref}},\dot{\omega}_{\text{ref}}]^{T}$, with each element constrained to the interval [1,1][-1,1]. These outputs are subsequently scaled by a factor of 0.5 for $\delta_{\text{ref}}$ and 5 for $\dot{\omega}_{\text{ref}}$ to yield the action vector. The steering scaling factor was chosen because 0.5 rad is the maximum limit of the steering actuator. The acceleration limit for the wheel was set to 5 $m/s^{2}$, as this was the highest longitudinal acceleration measured from the identification dataset. The reference steering angle $\delta_{\text{ref}}$ is directly applied to the servo for steering actuation, while the reference wheel speed derivative $\dot{\omega}_{\text{ref}}$ is integrated over time, analogously to [29], before being passed to the wheel speed controller. This formulation of speed control mitigates abrupt velocity changes, thereby enhancing policy transferability to the real world.
To effectively train racing policies, we designed the reward function to quantify the vehicle’s progress along the track and penalize boundary violations, and formulated it as follows:
where $\mathfrak{s}_{t}$ denotes the progress along the centerline expressed in Frenet coordinates [30]. Although the primary goal is to achieve faster laps, designing the reward function directly around this metric would hinder learning due to its sparse nature. Therefore, progress along the centerline is used as an effective proxy that provides denser feedback. This is why such a reward formulation, or similar variants, is commonly employed in other RL racing problems [3, 4, 5].
@@ IV Experiments
To determine the key factors that affect learning RL policies for autonomous racing, we conducted a series of experiments. A major focus of our study was to address the sim-to-real gap, for which we evaluated various domain randomization strategies (see Section IV-A) and the impact of actuator modeling (see Section IV-D) to identify the most effective approach to improve real-world transferability.
@@ IV-A Domain randomization
To evaluate the impact of domain randomization on both policy performance and safety, we trained networks using varying levels of parameter perturbation, specifically with $\sigma\in\{0.0,\,0.02,\,0.05,\,0.1\}$. Two randomization strategies were considered: (i) randomizing only the tire-track friction, acknowledging that friction can vary due to small changes in temperature and amount of dust on the track, and (ii) randomizing all single-track parameters (vehicle mass, inertia, weight distribution, etc.) to increase the policy’s robustness against model mismatch.
Policy performance was assessed by comparing both lap times and rack boundaries violations, which was quantified by the time-integral of the off-track distance, averaged per lap, measured in $m\cdot s]$. The off-track distance is defined as
@@
Figure 3: Real-world performance of networks trained with randomized friction and single-track parameters. Top: Lap times achieved with different types and $\sigma_{dr}$ of randomization. Bottom: $E_{off}$ and percentage of laps with crashes for different randomization configurations.
The results of the experiment with randomized friction levels and single track parameters are presented in Figure 3. As friction randomization increases, lap times tend to increase, while the integrated off-track error decreases. A similar effect was observed when randomizing all single-track parameters; however, overall, single-track parameter randomization caused worse performance, resulting in longer lap times and higher $E_{\text{off}}$ compared to policies trained with only friction randomization. This suggests that under more significant parameter variability, the policy adopts a more conservative driving strategy to remain valid across the entire distribution of dynamic models encountered during training. While domain randomization enhances robustness, excessive perturbations across all parameters can degrade overall performance. When the policy is trained over a wide distribution of models, it is forced to learn a control strategy that generalizes across the entire range of dynamics rather than optimizing for a specific one. Since the policy has no access to action history or recurrent memory to infer the underlying system dynamics at a given moment, it cannot adapt its behavior to a particular model instance. Instead, it must adopt a more cautious, suboptimal racing strategy that remains feasible across all possible parameter variations, ultimately leading to longer lap times.
These findings highlight the necessity of balancing domain randomization to ensure both safety and real-world transferability. Even with precise system identification, a moderate level of randomization is beneficial for mitigating simulation-to-reality discrepancies and improving policy robustness. However, while limited domain randomization improves generalization, high-performance policies still require accurate system identification, as randomization alone cannot fully compensate for an inaccurate dynamics model. For subsequent experiments, we selected the policy trained with friction randomization at $\sigma_{dr}=0.02$, and we refere to it as the base policy. This policy was selected for its acceptable trade-off between speed and robustness.